OpenStudy (anonymous):

Simplify the following expressions: 3(2x + 3 - x squared) - 2(2x squared - x- 3)

OpenStudy (anonymous):

first, simplify the terms \[3(2x + 3 -x^2) = 6x+9-3x^2\]and\[-2(2x^2 -x -3)= -4x^2 +2x +6\]

OpenStudy (anonymous):

where do you get the = sign??

OpenStudy (anonymous):

the left side is the original expression, on the right side the multiplied one I could also have used "gets" or something

OpenStudy (anonymous):

in total the multiplied expression looks like this \[6x+9-3x^2 -4x^2 +2x +6\]

OpenStudy (anonymous):

now i understand ;)

OpenStudy (anonymous):

so we need to collect the powers if they have the same base??

OpenStudy (anonymous):

YEP

OpenStudy (anonymous):

still not sure of my answers.

OpenStudy (anonymous):

\[8x+15-12x^{4}\] i think this solution is wrong.

OpenStudy (anonymous):

where did you get \( -12x^4\)? \(-4x^2-3x^2=?\) is not equal to \( -12x^4\) because you're adding them

OpenStudy (anonymous):

i multiply the base and did addition on the two exponent.

OpenStudy (anonymous):

year i now it wrong... i tried everything but still my answers are all wrong....can you please show me your working?? coz need to record it for me test next week.

OpenStudy (anonymous):

\(-4x^3-3x^2\) is not the same as \((-4x^2)(-3x^2)\) remember that when adding or subtracting variables, the exponents will be the same BUT if you are multiplying or dividing the exponents will change

OpenStudy (anonymous):

ok so it means that my answer is suppose to be -7xsquared +8x+15??

OpenStudy (anonymous):

yep

OpenStudy (mathmale):

phanta_seea, in her first post, typed out these two expressions:\[6x+9-3x^2\]\[-4x^2 +2x +6\]. All we need to do is to combine "like" terms. So, we'll arrange the "like" terms in columns: - 3x^2 + 6x + 9 - 4x^2 + 2x +6 ---------------- -7x^2 + 8x +15

OpenStudy (mathmale):

This is the shortest way to do this problem.

OpenStudy (anonymous):

thanks quys.